OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. m = 7, point (-2, 0)

jimthompson5910 (jim_thompson5910):

use y-y1 = m(x-x1) where in this case m = 7 x1 = -2 y1 = 0 plug that info into the equation, then solve for y if necessary

OpenStudy (anonymous):

y = 2x + 6

jimthompson5910 (jim_thompson5910):

y = 2x + 6 has a slope of ____

OpenStudy (anonymous):

i dont know:(

jimthompson5910 (jim_thompson5910):

y = mx+b has a slope of m right?

OpenStudy (anonymous):

yeah:(

jimthompson5910 (jim_thompson5910):

y = 2x+6 what is the value of m?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

i added the 2 and 6

jimthompson5910 (jim_thompson5910):

y = mx+b y = 2x+6 notice how the m and the 2 match up so m = 2

OpenStudy (anonymous):

oh.... makes sense

OpenStudy (anonymous):

thank you !!!!!

jimthompson5910 (jim_thompson5910):

so y = 2x + 6 has a slope of 2, NOT 7 that's why y = 2x+6 cannot be the answer